✓ Board-verified
Mathematics · 2022 · 2 marks
CBSE 2022 · Region 5 · Set 1 · Q6
(a) Events A and B are such that P ( A )=1/2, P ( B )=7/12 and P (Ā ∪ B̄)=1/4 Find whether the events A and B are independent or not.
(a)
Events A and B are such that \[\mathrm{P}(\mathrm{~A})=\frac{1}{2}, \mathrm{P}(\mathrm{~B})=\frac{7}{12} \text { and } \mathrm{P}(\overline{\mathrm{~A}} \cup \overline{\mathrm{~B}})=\frac{1}{4} \] Find whether the events $\displaystyle A$ and $\displaystyle \mathrm{B}$ are independent or not.
OR
(b)
A box $\displaystyle \mathrm{B}_{1}$ contains $\displaystyle 1$ white ball and $\displaystyle 3$ red balls. Another box $\displaystyle \mathrm{B}_{2}$ contains $\displaystyle 2$ white balls and $\displaystyle 3$ red balls. If one ball is drawn at random from each of the boxes $\displaystyle \mathrm{B}_{1}$ and $\displaystyle \mathrm{B}_{2}$, then find the probability that the two balls drawn are of the same colour.
Marking-scheme solution
(a)
$\displaystyle P(A) = \dfrac{1}{2},\ P(B) = \dfrac{7}{12},\ P(\bar{A} \cup \bar{B}) = \dfrac{1}{4}$
\[P(\bar{A} \cup \bar{B}) = P(\overline{A \cap B}) = 1 - P(A \cap B)\]
\[P(A \cap B) = 1 - \frac{1}{4} = \frac{3}{4}\]
\[P(A) \times P(B) = \frac{1}{2} \times \frac{7}{12} = \frac{7}{24}\]
\[P(A) \times P(B) \neq P(A \cap B)\]
$\displaystyle \therefore$ $\displaystyle A$ and $\displaystyle B$ are not independent
Or(b) $\displaystyle P$ (both balls drawn are of same colour)
\[= P \text{ (both white) } + P \text{ (both red) }\]
\[= \frac{1}{4} \times \frac{2}{5} + \frac{3}{4} \times \frac{3}{5} = \frac{11}{20}\]
More from Probability
- Assertion: Two coins are tossed simultaneously. The probability of getting two heads, if it is known that at least one head comes up, is 1/3. Reason…2023
- If A and B are two events such that P(B)=1/5, P(A ∣ B)=2/3 and P(A ∪ B)=3/5, then P(A) is:2025
- A shop selling electronic items sells smartphones of only three reputed companies A, B and C because chances of their manufacturing a defective…2025
- Three friends A, B and C got their photograph clicked. Find the probability that B is standing at the central position, given that A is standing at…2022
- A box has 4 green, 8 blue and 3 red pens. A student picks up a pen at random, checks its colour and replaces it in the box. He repeats this process 3…2025
- Smoking increases the risk of lung problems. A study revealed that 170 in 1000 males who smoke develop lung complications, while 120 out of 1000…2026
- Let M and N be two events such that P(M)=0.6, P(N)=0.2 and P(M ∩ N)=0.5, then P(M^′ / N^′) is2025
- The probability distribution of a random variable X is: where k is some unknown constant. The probability that the random variable X takes the value…2024
CBSE Class 12 Mathematics past-paper question from the 2022 board exam, with the answer as CBSE’s own marking scheme gives it. Where our answers come from.